# On the classification of centralizers of ODOs:

### v1

@@ 3.2 Planar spectral curves
Observe that $S_{0}(P-\lambda,Q-\mu)$ is a squared matrix of size $n+m$ and entries in $\Sigma[\lambda,\mu]$. See [44], for the definition and main properties of differential resultants.
It was first proved by E. Previato in [38] that $h(\lambda,\mu)$ is a constant coefficient polynomial; see also [51]. For differential operators with coefficients in an arbitrary differential field, these results were proved in [44], Theorem 5.4, see also [32].
@@ 3.3 Space spectral curves
4 Spectral sheaves
To a commutative subalgebra $\mathcal{A}$ of the ring of differential operators $\Sigma[\partial]$ we associate an algebraic curve $\Gamma_{\mathcal{A}}$ and a vector bundle $\mathcal{F}_{\mathcal{A}}$ on $\Gamma_{\mathcal{A}}$, the so called spectral sheaf [51, 4, 41]. To achieve an effective computation of $\mathcal{F}_{\mathcal{A}}$, we look at this problem as a factorization problem of $L-\lambda$ over a new coefficient field $\Sigma(\Gamma_{\mathcal{A}})$ that contains the coefficient field $\Sigma$ of $L$ and the field of the curve $\mathbf{C}(\Gamma_{\mathcal{A}})$. Let us consider the ring
$\Sigma[\Gamma_{\mathcal{A}}]=\frac{\Sigma[\lambda,\overline{\mu}]}{[\texttt{BC}(\mathcal{A})]}.$ (15)
@@ 4 Spectral sheaves
This conjecture was proved in [33] for $\mathcal{A}=\mathcal{Z}(L)$ the centralizer of a Schrödinger operator $L$. We summarize in the following the achievements for planar and space spectral curves.
@@ 4.1 Factorization over planar curves
Given commuting differential operators $P$, $Q$ in $\Sigma[\partial]$ and the irreducible polynomial $f\in\mathbf{C}[\lambda,\mu]$ that generates the BC ideal $\texttt{BC}(P,Q)$. The differential ideal $[f]$ generated by $f$ in $\Sigma[\lambda,\mu]$ is a prime differential ideal [44, Theorem 7], and we can consider the differential domain
@@
The derivation $\partial$ in $\Sigma$, is extended to $\Sigma[\lambda,\mu]$, setting $\partial(\lambda)=\partial(\mu)=0$. A derivation $\tilde{\partial}$ is naturally defined on $\Sigma[\Gamma_{P,Q}]$ by ~$\tilde{\partial}(q+[f])=\partial(q)+[f]$, for $q\in\Sigma[\lambda,\mu]$, we denote $\tilde{\partial}$ by $\partial$ if there is no room for confusion. We denote by $\Sigma(\Gamma_{P,Q})$ the fraction field of $\Sigma[\Gamma_{P,Q}]$, which is a differential field with the extended derivation.
The differential resultant of $P-\lambda$ and $Q-\mu$ is zero in $\Sigma(\Gamma_{P,Q})$, by the differential resultant theorem, see for instance [44, Theorem A5], then $P-\lambda$ and $Q-\mu$ have a nontrivial greatest common right divisor, as differential operators with coefficients in $\Sigma(\Gamma_{P,Q})$
@@
which is a global factor in the following sense. For every $(\lambda_{0},\mu_{0})\in\Gamma_{P,Q}\backslash Z$, where $Z$ is a finite subset of $\mathbf{C}^{2}$, we obtain a right factor in $\Sigma[\partial]$ of
@@
$\rm ord(\mathcal{F}(\lambda_{0},\mu_{0}))\leq\rm ord(\mathcal{F}(\lambda,\mu)),(\lambda_{0},\mu_{0})\in\Gamma_{P,Q}\backslash Z.$
@@
Theorem 4.1 (Wilson).
Let $\Sigma=\mathbf{C}((x))$ and consider commuting differential operators $P$, $Q$ in $\Sigma[\partial]$. Then the differential resultant
@@
For a rigorous proof see [51, Appendix]
@@
Let $\Sigma=\mathbf{C}((x))$. Given commuting differential operators $P$, $Q$ in $\Sigma[\partial]$, let us assume that ${\rm rank}(\mathbf{C}[P,Q])=r$. Then the greatest common right divisor of $P-\lambda$ and $Q-\mu$ as differential operators in $\Sigma(\Gamma_{P,Q})[\partial]$ has order $r$.
@@
Observe that $S_{k}$ is a matrix with $n+m-2k$ rows, $n+m-k$ columns and entries in $\Sigma[\lambda,\mu]$. For $i=0,\dots,k$ let $S_{k}^{i}$ be the squared matrix of size $n+m-2k$ obtained by removing the columns of $S_{k}$ indexed by $\partial^{k},\ldots,\partial,1$, except for the column indexed by $\partial^{i}$. The subresultant sequence of $P-\lambda$ and $Q-\mu$ is the next sequence of differential operators in $\Sigma[\lambda,\mu][\partial]$:
@@
By the subresultant theorem [6, Theorem 4], we obtain a closed form of the greatest common right divisor of $P-\lambda$ and $Q-\mu$ that can be effectively computed. Hence an explicit description of the rank $r$ spectral sheave $\mathcal{F}$ in the terminology of [51, 4, 41].
@@
Let $\Sigma=\mathbf{C}((x))$. Given commuting differential operators $P$, $Q$ in $\Sigma[\partial]$, let us assume that ${\rm rank}(\mathbf{C}[P,Q])=r$. Then the subresultants ${\rm s}\partial{\rm Res}_{k}(P-\lambda,Q-\mu)$ are identically zero in $\Sigma(\Gamma_{P,Q})[\partial]$, for $k=0,\ldots,r-1$ and
@@
$Z=\{(\lambda_{0},\mu_{0})\in\mathbf{C}^{2}\mid\phi_{r}(\lambda_{0},\mu_{0})=0\}.$ (18)
@@
where $\delta=x\partial$, is proved in [26] to be $\mathcal{Z}(L)=\mathbf{C}[L,B]$, where
@@
The second differential resultant equals
$\mathcal{F}(\lambda,\mu)=(\lambda x^{4}-560)\left(\frac{-\mu x^{2}+20\lambda}{x^{6}}\partial^{2}+\frac{-5(3\lambda x^{4}-1232)}{x^{9}}\partial+\frac{\lambda x^{4}-560}{x^{8}}\right).$
Observe that $\mathcal{F}(\lambda_{0},\mu_{0})$ is an order $2$ operator for every $(\lambda_{0},\mu_{0})\neq(0,0)$, the singular point.
Let us go back to differential operators with coefficients in an arbitrary differential field $\Sigma$. If ${\rm rank}(\mathbf{C}[P,Q])=1$ then it can be proved as [44, Lemma 8] that $\mathcal{F}$ is the order one differential operator determined by the first differential subresultant ${\rm s}\partial{\rm Res}_{1}(P-\lambda,Q-\mu)=\phi_{0}+\phi_{1}\partial$.
@@
By the construction of the matrices $S_{1}^{i}$ as in (17), the degree in $\lambda$ of $f$ is $\rm ord(Q)=m$ and the degree in $\mu$ is $\rm ord(P)=n$. By the construction of the matrices $S_{1}^{i}$ in (17), the degree in $\mu$ of $\phi_{i}(\lambda,\mu)$ is less than or equal to $n-1$. Therefore $\phi_{i}$ does not belong to $[\texttt{BC}(P,Q)]=[f]$. ∎
@@
By Lemma 4.3, ${\rm s}\partial{\rm Res}_{1}(P-\lambda,Q-\mu)$ is an order one operator and by the first subresultant theorem [44, Theorem B.4] it is the greatest common divisor $\mathcal{F}$ of $P-\lambda$ and $Q-\mu$. ∎
In some cases centralizers are isomorphic to coordinate rings of planar spectral curves. For instance, if $L$ is a differential operator of order $2$ then $\mathcal{A}=\mathcal{Z}(L)=\mathbf{C}[L,A_{1}]$ with $\rm ord(A_{1})\equiv_{2}1$. See also the family of Euler operators in Example 4.2.
@@ 4.2 Factorization over space curves
Let us assume that $\mathcal{A}=\mathcal{Z}(L)$ for an operator $L$ of order $3$ and whose basis as a $\mathbf{C}[L]$-module is $\{1,A_{1},A_{2}\}$. Then $\texttt{BC}(\mathcal{A})=(f_{1},f_{2},f_{3})$ and by [44, Theorem 11] the ideal $[\texttt{BC}(\mathcal{A})]$ generated by $\texttt{BC}(\mathcal{A})$ in $\Sigma[\lambda,\mu_{1},\mu_{2}]$ is a prime differential ideal. This allows to define the differential domain
@@
with the natural derivation and extend this derivation to its fraction field, denoted by $\Sigma(\Gamma_{\mathcal{A}})$.
@@
Let us assume that $\mathcal{A}=\mathcal{Z}(L)$ for an operator $L$ of prime order $p$ and whose basis as a $\mathbf{C}[L]$-module is $\{1,A_{1},A_{2}\}$. Then the monic greatest common right divisor $\mathcal{F}$ over $\Sigma(\Gamma_{\mathcal{A}})$ of $L-\lambda$, $A_{1}-\mu_{1}$ and $A_{2}-\mu_{2}$, equals the monic $\rm gcrd(L-\lambda,A_{i}-\mu_{i})$, $i=1,2$.
@@
In this situation [44, Proposition 1 and Lemma 9] are also satisfied. The proof of [44, Theorem 12] also applies to conclude that
@@
Consider the differential field $\Sigma=\mathbf{C}\langle\cosh(x)\rangle$ with derivation $\partial=d/dx$.
@@
$A_{1}=$ $\frac{16\eta^{2}-24}{\eta^{4}}+\frac{-24\eta^{\prime}}{\eta^{3}}\partial+\frac{(-4/3)\eta^{2}+8}{\eta^{2}}\partial^{2}+\partial^{4}$
@@
$+\frac{-40\eta^{\prime}}{\eta^{3}}\partial^{2}+\frac{10}{\eta^{2}}\partial^{3}+\partial^{5}$
@@
of $\Gamma_{\mathcal{A}}$ in ${\rm s}\partial{\rm Res}(L-\lambda,A_{i}-\mu_{i})$, $i=1,2$. The common factor $\mathcal{F}$ defines the spectral sheave on the spectral curve $\Gamma_{\mathcal{A}}$.
@@ 5 Spectral Picard-Vessiot fields
Following Drach’s ideology [12], we aim to compute the minimal field containing the solutions of the spectral problem $(L-\lambda)(y)=0$. Consider the field of the curve, the fraction field $\mathbf{C}(\Gamma_{\mathcal{A}})$ of the domain
@@
It is important to note that $\mathbf{C}(\Gamma_{\mathcal{A}})$ is not an algebraically closed field, which is required to prove existence of classical Picard-Vessiot fields [42]. Differential Galois theories do exist in the case of non algebraically closed field of constants, [CH], [27] but, to fully exploit its special properties, we aim to develop a Picard-Vessiot theory adapted to spectral problems $L(y)=\lambda y$, for a generic algebraic parameter $\lambda$ over the field of constants $\mathbf{C}$.
Let $\mathcal{A}=\mathcal{Z}(L)$, the centralizer of a Schrödinger operator $L=-\partial^{2}+u$, $u\in\Sigma$. Denote by $\partial-\phi_{+}$, resp. $\partial-\phi_{-}$, the monic greatest common right divisor of $L-\lambda$ and $A_{1}-\mu$, resp. $A_{1}+\mu$. Let $\Psi_{+}$, resp. $\Psi_{-}$, be a non zero solution of $(L-\lambda)(y)=0$ defined by
$\partial(\Psi_{+})=\phi_{+}\Psi_{+},\mbox{ resp. }\partial(\Psi_{-})=\phi_{-}\Psi_{-}.$
The solutions $\Psi_{+}$ and $\Psi_{-}$ belong to the differential closure of $\Sigma(\Gamma_{\mathcal{A}})$ and by [33, Lemma 4.5] are a fundamental system of solutions of $(L-\lambda)(y)=0$.
@@
    The field of constants of $(\Sigma(\Gamma_{\mathcal{A}}),\tilde{\partial})$ is $\mathbf{C}(\Gamma_{\mathcal{A}})$.
@@
For $\rm ord(L)>3$, it remains to prove that the differential ideal $[\texttt{BC}(\mathcal{A})]$ is a prime ideal in $\Sigma[\lambda,\overline{\mu}]$, before defining the new coefficient field $\Sigma(\Gamma_{\mathcal{A}})$ and computing the right factors of $L-\lambda$ over $\Sigma(\Gamma_{\mathcal{A}})$. Our aim is to prove the existence of spectral Picard-Vessiot fields for ODOs with non-trivial centralizers of rank $1$.
Determining if a commutative algebra $\mathcal{A}$ of ODOs has rank $1$ is a non trivial open problem to which differential Galois theory brings a new perspective. The differential Galois group of an ODO $L-\lambda$ was defined in [2], in the case of smooth spectral curves, and the centralizer of $L$ was proved to have rank $1$ if the group is commutative.
Acknowledgements. Supported by the grant PID2021-124473NB-I00, “Algorithmic Differential Algebra and Integrability” (ADAI) from the Spanish MICINN. The author is grateful to M.A. Zurro for numerous stimulating discussions and to A. Jiménez Pastor for computations of centralizers with the dalgebra package in SageMath.

### v2

@@ 3.2 Planar spectral curves
Observe that $S_{0}(P-\lambda,Q-\mu)$ is a squared matrix of size $n+m$ and entries in $\Sigma[\lambda,\mu]$. For the definition and main properties of differential resultants for ODOs see [10], [35] or [53].
It was first proved by G. Wilson in [61] that $h(\lambda,\mu)$ is a constant coefficient polynomial; see also [47]. For differential operators with coefficients in an arbitrary differential field, these results were proved in [63], Section 5.3, see also [53], Theorem 5.4 and [41].
@@ 3.3 Space spectral curves
4 Factorization over spectral curves
Given a commutative subalgebra $\mathcal{A}$ of the ring of differential operators $\Sigma[\partial]$, we review next how to define a vector bundle $\mathcal{F}_{\mathcal{A}}$ on the algebraic curve $\Gamma_{\mathcal{A}}$. From the extense existing literature, our point of view is closer to the results of Previato and Wilson [61, 50].
In order to achieve an effective computation of $\mathcal{F}_{\mathcal{A}}$, we look at this problem as a factorization problem of $L-\lambda$ over a new coefficient field $\Sigma(\Gamma_{\mathcal{A}})$, that contains the coefficient field $\Sigma$ of $L$ and the field of the curve $\mathbf{C}(\Gamma_{\mathcal{A}})$. Let us consider the ring
$\Sigma[\Gamma_{\mathcal{A}}]:=\frac{\Sigma[\lambda,\overline{\mu}]}{[\texttt{BC}(\mathcal{A})]}.$ (15)
@@ 4 Factorization over spectral curves
This conjecture was proved for $\mathcal{A}=\mathcal{Z}(L)$, for a Schrödinger operator $L$, in [42]. We summarize in the following the achievements for planar and space spectral curves in [53]. It would be interesting to clarify the relation between the so called spectral sheaf on the points of the spectral curve [6] and the vector bundles that we will compute by means of greatest common right divisors in $\Sigma(\Gamma_{\mathcal{A}})[\partial]$.
@@ 4.1 Factorization over planar curves
Given commuting differential operators $P$, $Q$ in $\Sigma[\partial]$ and the irreducible polynomial $f\in\mathbf{C}[\lambda,\mu]$ that generates the BC ideal $\texttt{BC}(P,Q)$, then the differential ideal $[f]$ generated by $f$ in $\Sigma[\lambda,\mu]$ is a prime differential ideal [53, Theorem 7], and we can consider the differential domain
@@
The derivation $\partial$ in $\Sigma$, is extended to $\Sigma[\lambda,\mu]$, setting $\partial(\lambda)=\partial(\mu)=0$. A derivation $\tilde{\partial}$ is naturally defined on $\Sigma[\Gamma_{P,Q}]$ by $\tilde{\partial}(q+[f])=\partial(q)+[f]$, for $q\in\Sigma[\lambda,\mu]$, we denote $\tilde{\partial}$ by $\partial$ if there is no room for confusion. We denote by $\Sigma(\Gamma_{P,Q})$ the fraction field of $\Sigma[\Gamma_{P,Q}]$, which is a differential field with the extended derivation.
The differential resultant of $P-\lambda$ and $Q-\mu$ is zero in $\Sigma(\Gamma_{P,Q})$, by the differential resultant theorem, see for instance [53, Theorem A5], then $P-\lambda$ and $Q-\mu$ have a nontrivial greatest common right divisor, as differential operators with coefficients in $\Sigma(\Gamma_{P,Q})$
@@
which is a global factor in the following sense. For every $\Z(\lambda_{0},\mu_{0})\in\Gamma_{P,Q}\backslash$ Z, where $Z$ is a finite subset of $\mathbf{C}^{2}$, we obtain a right factor in $\Sigma[\partial]$ of
@@
$\Z.\rm ord(\mathcal{F}(\lambda_{0},\mu_{0}))\leq\rm ord(\mathcal{F}(\lambda,\mu)),(\lambda_{0},\mu_{0})\in\Gamma_{P,Q}\backslash$ Z.
@@
Theorem 4.1 ( G. Wilson[61]).
Let $\Sigma=\mathbb{C}((x))$ and consider commuting differential operators $P$, $Q$ in $\Sigma[\partial]$. Then the differential resultant
@@
For a rigorous proof see [61, Appendix]
@@
Let $\Sigma=\mathbb{C}((x))$. Given commuting differential operators $P$, $Q$ in $\Sigma[\partial]$, let us assume that ${\rm rank}(\mathbf{C}[P,Q])=r$. Then the greatest common right divisor of $P-\lambda$ and $Q-\mu$ as differential operators in $\Sigma(\Gamma_{P,Q})[\partial]$ has order $r$.
@@
Observe that $S_{k}$ is a matrix with n+m2kn+m-2k rows, n+mkn+m-k columns and entries in $\Sigma[\lambda,\mu]$. For $i=0,\dots,k$ let $S_{k}^{i}$ be the squared matrix of size n+m2kn+m-2k obtained by removing the columns of $S_{k}$ indexed by $\partial^{k},\ldots,\partial,1$, except for the column indexed by $\partial^{i}$. The subresultant sequence of $P-\lambda$ and $Q-\mu$ is the next sequence of differential operators in $\Sigma[\lambda,\mu][\partial]$:
@@
By the subresultant theorem [10, Theorem 4], we obtain a closed form of the greatest common right divisor of $P-\lambda$ and $Q-\mu$ that can be effectively computed. Hence an explicit description of the rank $r$ vector bundle $\mathcal{F}$ over $\Gamma_{P,Q}$ is obtained [50].
@@
Let $\Sigma=\mathbb{C}((x))$. Given commuting differential operators $P$, $Q$ in $\Sigma[\partial]$, let us assume that ${\rm rank}(\mathbf{C}[P,Q])=r$. Then the subresultants ${\rm s}\partial{\rm Res}_{k}(P-\lambda,Q-\mu)$ are identically zero in $\Sigma(\Gamma_{P,Q})[\partial]$, for $k=0,\ldots,r-1$ and
@@
$Z=\{(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}\mid\phi_{r}(\lambda_{0},\mu_{0})=0\}.$ (18)
We conjecture that $\mathcal{F}$ is irreducible as a differential operator with coefficients in $\Sigma(\Gamma_{P,Q})$, that is for generic values of $\lambda$ and $\mu$, and only for a finite set of points on $\Gamma_{P,Q}$ it is reducible. The next example illustrates Corollary 4.1.2 and this conjecture.
@@
where $\delta=x\partial$, is proved in [34] to be $\mathcal{Z}(L)=\mathbb{C}[L,B]$, where
@@
The spectral curve $\Gamma_{L,B}$ of the pair $L,B$ is thus defined by $\lambda^{3}-\mu^{2}$. We can set $\Sigma=\mathbb{C}(x)$ and consider the new coefficient field $\Sigma(\Gamma_{L,B})$, which is the quotient field of
$\Sigma[\Gamma_{L,B}]=\frac{\Sigma[\lambda,\mu]}{[\lambda^{3}-\mu^{2}]}.$
@@
The second differential subresultant equals
$\mathcal{F}(\lambda,\mu)=(\lambda x^{4}-560)\left(\frac{-\mu x^{2}+20\lambda}{x^{6}}+\frac{-5(3\lambda x^{4}-1232)}{x^{9}}\partial+\frac{\lambda x^{4}-560}{x^{8}}\partial^{2}\right).$
Observe that $\phi_{2}(\lambda,\mu)=\frac{(\lambda x^{4}-560)^{2}}{x^{8}}$ is non zero for every $(\lambda,\mu)$ in $\mathbb{C}^{2}$. The first differential subresultant equals $(\lambda x^{4}-560)(\lambda^{3}-\mu^{2})/x^{4}$, which is identically zero in $\Sigma(\Gamma_{P,Q})[\partial]$.
At the singular point $\mathcal{F}(0,0)=(\partial-\frac{11}{x})\partial$, while $\mathcal{F}(1,1)$ is irreducible. We conjecture that $\mathcal{F}(\lambda_{0},\mu_{0})$ will be irreducible for every non singular $\lambda_{0},\mu_{0}$ of $\Gamma_{L,B}$.
Let us go back to differential operators with coefficients in an arbitrary differential field $\Sigma$. If ${\rm rank}(\mathbf{C}[P,Q])=1$ then it can be proved as [53, Lemma 8] that $\mathcal{F}$ is the order one differential operator determined by the first differential subresultant ${\rm s}\partial{\rm Res}_{1}(P-\lambda,Q-\mu)=\phi_{0}+\phi_{1}\partial$.
@@
By the construction of the matrices $S_{1}^{i}$ as in (17), the degree in $\lambda$ of $f$ is $\rm ord(Q)=m$ and the degree in $\mu$ is $\rm ord(P)=n$. By the construction of the matrices $S_{1}^{i}$ in (17), the degree in $\mu$ of $\phi_{i}(\lambda,\mu)$ is less than or equal to n1n-1. Therefore $\phi_{i}$ does not belong to $\texttt{BC}(P,Q)]=[f]$. ∎
@@
By Lemma 4.3, ${\rm s}\partial{\rm Res}_{1}(P-\lambda,Q-\mu)$ is an order one operator and by the first subresultant theorem [53, Theorem B.4] it is the greatest common divisor $\mathcal{F}$ of $P-\lambda$ and $Q-\mu$. ∎
In some cases centralizers are isomorphic to coordinate rings of planar spectral curves. For instance, if $L$ is a differential operator of order $2$ then $\mathcal{A}=\mathcal{Z}(L)=\mathbf{C}[L,A]$ with $\rm ord(A)\equiv_{2}1$. See also the family of Euler operators in Example 4.2.
@@ 4.2 Factorization over space curves
Let us assume that $\mathcal{A}=\mathcal{Z}(L)$ for an operator $L$ of order $3$ and whose basis as a $\mathbf{C}[L]$-module is $\{1,A_{1},A_{2}\}$. Then by [53, Theorem 10] $\texttt{BC}(\mathcal{A})=(f_{1},f_{2},f_{3})$ and by [53, Theorem 11] the ideal $\texttt{BC}(\mathcal{A})]$ generated by $\texttt{BC}(\mathcal{A})$ in $\Sigma[\lambda,\mu_{1},\mu_{2}]$ is a prime differential ideal. This allows to define the differential domain
@@
with the natural derivation and extend this derivation to its fraction field $\Sigma(\Gamma_{\mathcal{A}})$.
@@
Let us assume that $\mathcal{A}=\mathcal{Z}(L)$ for an operator $L$ of order $3$ whose basis as a $\mathbf{C}[L]$-module is $\{1,A_{1},A_{2}\}$. Then the monic greatest common right divisor $\mathcal{F}$ over $\Sigma(\Gamma_{\mathcal{A}})$ of $L-\lambda$, $A_{1}-\mu_{1}$ and $A_{2}-\mu_{2}$, equals the monic $\rm gcrd(L-\lambda,A_{i}-\mu_{i})$, $i=1,2$.
@@
In this situation [53, Proposition 1 and Lemma 9] are satisfied and [53, Theorem 12] applies to conclude that
@@
Consider the differential field $\Sigma=\mathbf{C}\langle\cosh(x)\rangle$ with derivation $\partial=d/dx$ and
@@
$A_{1}=$ $\frac{16\eta^{2}-24}{\eta^{4}}+\frac{-24\eta^{\prime}}{\eta^{3}}\partial+\frac{(-4/3)\eta^{2}+8}{\eta^{2}}\partial^{2}+\partial^{4},$
@@
$+\frac{-40\eta^{\prime}}{\eta^{3}}\partial^{2}+\frac{10}{\eta^{2}}\partial^{3}+\partial^{5}.$
@@
of $\Gamma_{\mathcal{A}}$ in ${\rm s}\partial{\rm Res}(L-\lambda,A_{i}-\mu_{i})$, $i=1,2$. The common factor $\mathcal{F}$ defines the line bundle on the spectral curve $\Gamma_{\mathcal{A}}$.
@@ 5 Spectral Picard-Vessiot fields
Following Drach’s ideology [15], we aim to compute the minimal field containing the solutions of the spectral problem $(L-\lambda)(y)=0$. Consider a finitely generated commutative $\mathbf{C}$-algebra $\mathcal{A}$ in $\Sigma[\partial]$ as in (9) and the field of the curve $\Gamma_{\mathcal{A}}$, the fraction field $\mathbf{C}(\Gamma_{\mathcal{A}})$ of the domain
@@
It is important to note that $\mathbf{C}(\Gamma_{\mathcal{A}})$ is not an algebraically closed field, which is required to prove existence of classical Picard-Vessiot fields [51]. Differential Galois theories do exist in the case of non algebraically closed field of constants, [11], [36] but, to fully exploit its special properties, we aim to develop a Picard-Vessiot theory adapted to spectral problems $L(y)=\lambda y$, for a generic algebraic parameter $\lambda$ over the field of constants $\mathbf{C}$.
Let $\mathcal{A}=\mathcal{Z}(L)=\mathbf{C}[L,A]$ be the centralizer of a Schrödinger operator $L=-\partial^{2}+u$, $u\in\Sigma$. Denote by $\partial-\phi_{+}$, resp. $\partial-\phi_{-}$, the monic greatest common right divisor of $L-\lambda$ and $A-\mu$, resp. $A+\mu$. Let $\Psi_{+}$, resp. $\Psi_{-}$, be a non zero solution of $(L-\lambda)(y)=0$ defined by
[…]
The solutions $\Psi_{+}$ and $\Psi_{-}$ belong to the differential closure of $\Sigma(\Gamma_{\mathcal{A}})$ and by [42, Lemma 4.5] are a fundamental system of solutions of $(L-\lambda)(y)=0$.
@@
[…]
@@
For $\rm ord(L)>3$, it remains to prove that the differential ideal $\texttt{BC}(\mathcal{A})]$ is a prime ideal in $\Sigma[\lambda,\overline{\mu}]$, before defining the new coefficient field $\Sigma(\Gamma_{\mathcal{A}})$ and computing the right factors of $L-\lambda$ over $\Sigma(\Gamma_{\mathcal{A}})$. Our aim is to prove the existence of spectral Picard-Vessiot fields for ODOs with non-trivial centralizers of rank $1$.
Determining if a commutative algebra $\mathcal{A}$ of ODOs has rank $1$ is a non trivial open problem to which differential Galois theory brings a new perspective. The differential Galois group of an ordinary differential operator $L-\lambda$ was defined in [3], and the centralizer of $L$ was proved to have rank $1$ if the group is commutative. More precisely in [3] they define the differential Galois group of a quantum complete integrable system (QCIS). It is important to note, that in the definition of differential Galois group given in [3] the parameter $\lambda$ is a generic parameter, a generic constant with respect to the derivation. It would be interesting to understand its relation with the differential Galois group of the parametric Picard-Vessiot theory introduced by Cassidy and Singer in [9], or appropriate extensions of this theory [18].
Other authors have studied the differential Galois group of spectral problems $Ly=\lambda_{0}y$, with $\lambda=\lambda_{0}$ an specialization of the spectral parameter [4, 20, 1], by means of the differential Galois group of the classical Picard-Vessiot theory, where the field of constants is algebraically closed. Our aim is to develop algorithms to compute differential Galois groups with parameters of spectral problems $Ly=\lambda y$, in the same spirit of the existing ones for differential Galois groups [56], [23], [17] or parametrized differential Galois groups [2], [36], to determine if the centralizer of $L$ has rank $1$.
Acknowledgments. Supported by the grant PID2021-124473NB-I00, “Algorithmic Differential Algebra and Integrability” (ADAI) from the Spanish MCIN/AEI /10.13039/501100011033 and by FEDER, UE. The author is grateful to M.A. Zurro for numerous stimulating discussions, to A. Jiménez Pastor for computations of centralizers with the dalgebra package in SageMath and to the anonymous referee whose remarks allowed to improve the exposition.